using addition principle solve
x+1/11=5/11 answer must be simplified. Type a fraction.
Thank you so very much.

Answer by stanbon(75887)

x+1/11=5/11
Your problem can be read in two ways:
---------------
1st:
x+(1/11) = 5/11
x = 4/11
----------------
2nd:
(x+1)/11 = 5/11
Multiply both sides by 11 to get:
x+1 = 5
x = 4
---------
It is important to use parentheses so others will read the problem
the way you see it.
------------
Cheers,
Stan H.